OpenStudy (anonymous):

Will give a medal and fan... Ralph is 27 years old and starting an IRA (individual retirement account). He is going to invest $200 at the beginning of each month. The account is expected to earn 2.65% interest, compounded monthly. How much money will Ricky have in his IRA when he retires, at age 65? $157,419.08 $94,723.10 $13,183.51 $416,424.15

OpenStudy (anonymous):

I'll try to help, hold on. :)

OpenStudy (anonymous):

Okay :D

OpenStudy (anonymous):

Hmmm...I don't get any of those answers listed...

OpenStudy (anonymous):

Do you know if there is a formula you can use?

OpenStudy (anonymous):

Let me look it up lol I am sorry it took me a moment to respond I was working on something else

OpenStudy (anonymous):

It's fine!

OpenStudy (anonymous):

A = P(1 + r)^t this is the compound interest formula

OpenStudy (anonymous):

FV = PV(1 + i)^nt this is the future value formula

OpenStudy (anonymous):

Ok, so, using those formulas, what do you get? :)

OpenStudy (anonymous):

I am not sure because I do not think this is the correct formula to solve this problem....

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

this is a loan of 0, with payments that are positive instead of negative; in other words B=0 and we are left with:\[A=P\frac{1-k^n}{1-k}\]

OpenStudy (amistre64):

that might be a negative so just ignore the sign

OpenStudy (anonymous):

Okay... what would I plug in for k? And P would be 200 correct?

OpenStudy (amistre64):

P = 200, k is the same compounding stuff as usual

OpenStudy (anonymous):

So .0265/12?

OpenStudy (amistre64):

1+ but yeah

OpenStudy (anonymous):

n= 12?

OpenStudy (amistre64):

the number of periods it is compounded is 27 to 65 years, 12 times a year

OpenStudy (amistre64):

oh, and since its a beginning of the month, we can assume an initial Balance of 200, not 0

OpenStudy (amistre64):

payments still add to it so we dont have to subtract any payments

OpenStudy (anonymous):

so we do not subtract 1 then?

OpenStudy (amistre64):

why would we subtract 1? im assuming your using the formulas i developed

OpenStudy (anonymous):

Yes I am lol

OpenStudy (anonymous):

Idk I am so confused... Im going to try to do it and see if I get it correct lol... one second

OpenStudy (amistre64):

\[A = Bk^n+P\frac{1-k^n}{1-k}\] we start with a beginning balance of 200, and payments of 200

OpenStudy (amistre64):

this format works well for loans to if we assume the B is negative, and we are trying to fill in the hole with payments

OpenStudy (amistre64):

*too not that spelling counts in math lol

OpenStudy (anonymous):

I got 35,395.01 0.o

OpenStudy (anonymous):

I did something wrong....

OpenStudy (amistre64):

prolly the time factor

OpenStudy (anonymous):

Yes it was the time that I did wrong... But I understand how to do it now... thank you so much! :D

OpenStudy (amistre64):

youre welcome

OpenStudy (amistre64):

im assuming the 619 is the result of me having an extra little 200 in the mix, but its close enough eh :)